OpenStudy (ericanoel912):

Algebra II help plz- WILL FAN & MEDAL Can someone walk me through how to find the inverse of a function?? Given the function f(x) = 6(x+2) − 3, solve for the inverse function when x = 21. −45 0 2 63

OpenStudy (imstuck):

To find the inverse function, the first thing you have to do is switch the x and y and solve for the new y. Then fill in the 21 for the new x. Like this:

OpenStudy (imstuck):

\[f(x) = y\]so\[y=6(x+2)-3\]Distribute and do the math to get\[y=6x+9\]

OpenStudy (imstuck):

now switch the x and the y and solve for the new y.

OpenStudy (imstuck):

\[x=6y+9\]Subtract 9 from both sides now to get\[x-9=6y\]and divide both sides by 6 to get the y alone.\[\frac{ x-9 }{ 6 }=y\]

OpenStudy (imstuck):

Now fill in the x with 21 and solve for y. What do you get when you do that?

OpenStudy (ericanoel912):

y=2!!

OpenStudy (ericanoel912):

Thank you so much! My Algebra II teacher couldn't explain it that well :) You're awesome!